# Enhanced many-body quantum scars from the non-Hermitian Fock skin effect

### v1

@@ Abstract
In contrast with extended Bloch waves, a single particle can become spatially localized due to the so-called skin effect originating from non-Hermitian pumping. Here we show that in a wide class of kinetically-constrained many-body systems, the skin effect can instead manifest as dynamical amplification within the Fock space, beyond the intuitively expected and previously studied particle localization and clustering. We exemplify this non-Hermitian Fock skin effect in an asymmetric version of the PXP model and show that it gives rise to ergodicity-breaking eigenstates – the non-Hermitian analogs of quantum many-body scars. A distinguishing feature of these non-Hermitian scars is their enhanced robustness against external disorders. We propose an experimental realization of the non-Hermitian scar enhancement in a tilted Bose-Hubbard optical lattice with laser-induced loss. Our results show that the Fock skin effect provides a powerful tool for creating robust non-ergodic states in generic open quantum systems.
@@
The experimental observations of QMBSs in versatile synthetic platforms ranging from ultracold atoms to superconducting circuits [11, 29, 30, 31, 32, 33, 34] have revealed the intrinsic fragility of QMBS dynamical revivals. In some cases, this has been mitigated by periodic driving, which can extend the lifetime of QMBS oscillations [29, 32, 35, 36]. Nevertheless, thermalization or decay to a steady state, due to inherent decoherence from atom loss or thermal fluctuations, ultimately sets in. Thus, achieving stable QMBSs in real-world settings, with unavoidable non-unitarity, remains a challenge for both theory and experiment [37, 38, 39, 40, 41].
In this paper, we show that non-Hermiticity presents a new route for enhancing ergodicity breaking and many-body scarring. Borrowing inspiration from well-known mechanisms such as non-Hermitian pumping [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], we demonstrate that a class of many-body systems with kinetic constraints can admit a form of “Fock skin effect” that amplifies the revivals of special initial states and makes them resilient against spatial disorder. Unlike the conventional non-Hermitian skin effect (NHSE), the Fock skin effect is characterized not by the uni-directional spatial evolution of wavefunctions, but by persistent dynamical recurrences in the Fock space of a many-body system. We demonstrate this mechanism of QMBS enhancement in a non-Hermitian variant of the PXP model, which can be realized in a Bose–Hubbard optical lattice with laser-induced loss [56, 57, 58, 59, 60]. Connection with conventional non-Hermitian pumping is made through the “forward scattering approximation” [12, 61, 22], which effectively maps the many-body system to a spatially-inhomogeneous tight-binding chain with non-reciprocal hopping amplitudes. Our results offer a new perspective on obtaining stable many-body revivals through dissipative perturbations, extending beyond recent approaches based on decoherence-free subspaces [62, 63].
Figure 1: (a) The action of operators in Eq. (1). Blue directed arrows label the asymmetric couplings $\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}$ (arrow for enhanced spin flipping $(1{+}u)\hat{X}^{+}_{j}$), while red links represent the symmetric coupling $\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$. (b) The directed Fock-space graph for the non-Hermitian PXP model in Eq. (1) with $L=6$ sites under PBCs. Each vertex represents a Fock state. The vertices are organized into vertical layers according to their Hamming distance $x$ from the state $|\ket{\mathbb{Z}_{2}}\equiv\ket{\uparrow\downarrow\uparrow\downarrow\ldots}$, located at the left edge of the graph. The translated Néel state, |¯$|\ket{\bar{\mathbb{Z}}_{2}}\equiv\ket{\downarrow\uparrow\downarrow\uparrow\ldots}$, is at the opposite end of the graph. The edges between Fock states represent the transitions allowed by the Hamiltonian in Eq. (1). Notably, all directed edges point towards $x{=}0$, denoting asymmetric pumping from the Fock skin effect.
@@
Here, $\hat{X}$, $\hat{Y}$, $\hat{Z}$ denote the standard Pauli matrices, such that $\hat{P}{=}(\hat{I}{-}\hat{Z})/2$ projects onto the spin-$\downarrow$ state. Non-Hermiticity enters through the asymmetric spin flip operator, $\hat{X}^{\prime}{=}\hat{X}{+}iu\hat{Y}$, acting on the even sites, FIG. 1(a), where $u$ controls the strength of non-Hermiticity. Each symmetric $\hat{X}$ or antisymmetric $\hat{X}^{\prime}$ spin flip acts only when its neighboring spins are both spin-$\downarrow$. To distinguish our Fock skin effect from the conventional skin effect marked by spatial boundary accumulation, we remove the physical boundaries by imposing periodic boundary conditions (PBCs). Moreover, we restrict to the $10{<}u{<}$ 1 regime, where the energy spectrum is real, circumventing wavefunction decay induced by complex spectra, see Supplementary Material (SM) [66].
@@
To provide insight behind the non-Hermitian scar enhancement, we analyze the wavefunction dynamics within the Fock space. The latter can be visualized as a quasi-1D network structure by grouping Fock states according to their Hamming distance from the Néel state, $|\ket{\mathbb{Z}_{2}}=\ket{\uparrow\downarrow\uparrow\downarrow\cdots}$, see an example in FIG. 1. Note that the |$\ket{\mathbb{Z}_{2}}$ state is the one that exhibits QMBS revivals in the Hermitian PXP model [12]. The Hamming distance indicates how many applications of the Hamiltonian are needed to reach the states in a given layer starting from the |$\ket{\mathbb{Z}_{2}}$ state. This is because the Hamiltonian only connects layer $x$ with $x\pm 1$, as shown by edges in FIG. 1(b), making apparent the connection with a tight-binding chain. Crucially, this picture reveals a consistent direction in the asymmetry of the transitions between Fock states. As illustrated, the blue directional transitions are all directed towards the |$\ket{\mathbb{Z}_{2}}$ state at $x{=}0$, reminiscent of a conventional NHSE chain but with Fock states instead of physical sites. However, with multiple Fock states in each layer instead of a single site, the “boundary” skin localization would be weak [66], and we instead expect more pronounced QMBS revivals, rather than the directional bulk current in conventional NHSE systems [43, 46, 67, 68].
@@
The stratification of the Fock space into Hamming layers can be formally expressed by generalizing the forward scattering approximation (FSA) [12, 61], which allowed to understand the origin of QMBSs in the Hermitian case [22]. We first decompose our non-Hermitian PXP Hamiltonian in Eq.(1) as $\hat{H}{=}\hat{H}^{+}{+}\hat{H}^{-}$, where
$\hat{H}^{+}=$ $\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime-}_{j}\hat{P}_{j+1}\!+\!\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}^{+}_{j}\hat{P}_{j+1},$ (2)
$\hat{H}^{-}=$ $\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime+}_{j}\hat{P}_{j+1}\!+\!\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}^{-}_{j}\hat{P}_{j+1},$
@@
with $\hat{X}^{\pm}{\equiv}(\hat{X}\pm i\hat{Y})/2$ and $\hat{X}^{\prime\pm}{\equiv}(1\pm u)(\hat{X}\pm i\hat{Y})/2$. By successively acting with $\hat{H}^{+}$ on the |$\ket{\mathbb{Z}_{2}}$ state, which we label as |$\ket{0}_{\rm FSA}$, we can generate a tower of $L+1$ FSA basis states |$\ket{n}_{\mathrm{FSA}}{\propto}(H^{+})^{n}\ket{0}_{\mathrm{FSA}}$, with $n=0,1,2,\ldots,L$ [61]. These states span the Fock-space layers in FIG. 1(b). In the FSA basis, the projected Hamiltonian (1) is given by
@@
with the non-uniform and asymmetric hopping amplitudes
$t^{+}_{x}\!=\!\bra{x\!+\!1}_{\rm FSA}\!\!\hat{H}^{+}\!\!\ket{x}_{\rm FSA},\quad t^{-}_{x}\!=\!\bra{x}_{\rm FSA}\!\!\hat{H}^{-}\!\!\ket{x\!+\!1}_{\rm FSA}\!.$ (4)
@@
Despite the formal analogy between the model (3) and a tight-binding chain, the FSA basis |$\ket{x}_{\mathrm{FSA}}$ represents the superposition of all Fock states spanning Hamming layer $x$. Moreover, both the basis states and the effective hopping amplitudes $t^{\pm}$ intricately depend on the non-hermiticity $u$ of the Hamiltonian, as we show below.
@@
Figure 2: (a) Normalized fidelity $F(t)$ for the initial state |$\ket{\mathbb{Z}_{2}}$ in the non-Hermitian PXP model (1). Turning on moderate non-Hermiticity $u\neq 0$ significantly enhances the return probability to the initial state. For large $u$ (dashed green curve), robust revivals are still present, but they become irregular due to multiple frequencies in the dynamics. (b) Amplitudes of effective hoppings in Eq.(4). The asymmetry between $t^{\pm}_{n}$ for $u{>}0$ (circles vs. triangles) leads to the non-Hermitian Fock skin effect in our model. (c) The concentration of eigenstates in the Hamming layer $x$, Eq. (5). A sharp Fock skin mode emerges under strong non-Hermiticity (green curve), corresponding to the persistent revivals in (a). (d) Comparison between the FSA left-boundary localization, $|\bra{\mathbb{Z}_{2}}\ket{\phi_{j}}_{\rm FSA}|$ |, and overlaps of |$\ket{\mathbb{Z}_{2}}$ with all eigenstates of the Hamiltonian in Eq.(1). From $u{=}0.0$ to $u{=}0.2$, the boundary localization from FSA vectors is in close agreement with the overlap of $\mathbb{Z}_{2}$ state with the equally-spaced QMBS states. All data is for a chain of size $L{=}16$ with PBCs.
@@
QMBS enhancement due to Fock skin effect.—QMBSs manifest as pronounced revivals in the fidelity, $F(t)\equiv|{\bra{\psi(0)}\ket{\psi(t)}}|^{2}/{\bra{\psi(t)}\ket{\psi(t)}}$, following a global quench from a chosen initial state, typically the |$\ket{\psi(0)}=\ket{\mathbb{Z}_{2}}$ state for PXP-type models [11, 12]. The time evolved state is obtained by |$\ket{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, and we explicitly normalize the fidelity at each time step. Strong revivals in $F(t)$ indicate broken ergodicity since the fidelity should decay rapidly under thermalizing dynamics [12, 13]. In FIG. 2(a), the fidelity $F(t)$ is plotted for the PXP model (1) with various degrees of non-Hermiticity $u$. The decay of peaks in $F(t)$ is noticeably slower for non-Hermitian $u{=}0.2$ (blue) as compared to the Hermitian $u{=}0$ case (red). With much larger non-Hermiticity $u{=}0.7$ (green dashed), both frequency and magnitude of the revivals are significantly enhanced, such that no significant decay is observed even after $t{=}10$.
@@
To explicitly attribute this non-Hermitian scar enhancement to the Fock skin effect, we plot the hopping amplitudes (4) in FIG. 2(b). Saliently, the FSA hoppings are asymmetric with $t_{x}^{+}$ (round markers) and $t_{x}^{-}$ (triangle markers) of different magnitudes, with the asymmetry becoming more pronounced with greater non-Hermiticity $u$. We note that the hopping asymmetry is markedly different from that in the physical Hamiltonian (1), since $x$ is the Hamming distance from the |$\ket{\mathbb{Z}_{2}}$ state in Fock space and not a spatial coordinate. Importantly, the asymmetric hopping amplitudes are non-trivially dependent on $x$ but asymmetric in the same direction ($t_{x}^{+}\geq t_{x}^{-}$) within the entire FSA chain. We emphasize that this property is not generic, see SM [66] for the discussion of a related model where this is not the case, with crucial implications on the existence of Fock skin accumulation.
@@
As seen in FIG. 2(c), the Fock skin pumping of $u{=}0.2$ shifts the symmetric $\Pi_{\rm Fock}(x)$ profile of $u{=}0$ slightly towards left, yet without pronounced “boundary” localization. Nevertheless, this increased proximity to the |$\ket{\mathbb{Z}_{2}}$ state is sufficient to noticeably enhance the scarred revivals [FIG. 2(a), blue]. A much stronger non-Hermiticity of $u{=}0.7$ results in a much more skewed $\Pi_{\rm Fock}(x)$, but still much less localized than conventional skin states corresponding to $t_{x}^{+}/t_{x}^{-}{\approx}3$ [FIG. 2 (b), green], with localization lengths of about $[\log 3]^{-1}{\approx}1$ site. Crucially, this lack of conventional exponential boundary skin localization allows for significant state diffusion in the Hilbert space, but with enhanced return probability to the |$\ket{\mathbb{Z}_{2}}$ state.
@@
Robustness to disorder.—We have established that the Fock skin effect can significantly enhance QMBSs. Given that the real-space skin effect can result in robust, directed dynamical currents, we posit that the non-Hermiticity in our system can not only amplify the quantum scar but also protect it against external (Hermitian) perturbations. To check this, we investigate the robustness of quantum scar under spatial disorder,
$\hat{H}_{\rm dis}=\hat{H}+\sum_{j}w_{j}\hat{Z}_{j},$ (6)
where $w_{j}$ denotes the strength of on-site random potential drawn from a uniform distribution within the range $[-W/2,W/2]$. It was previously shown that Hermitian QMBS revivals decay with increasing disorder strength [70]. For our model with the Fock skin effect at up to $u{=}0.35$, this resilience under moderate disorders is remarkably highlighted by the fidelity shown in FIG. 3(a), where robust scarring signatures are still present, in contrast to the rapid collapse at Hermitian $u{=}0$ without the Fock skin effect. The quantitative demonstration of enhanced stability is illustrated through the first fidelity peak, $F_{1}={\rm Max}[F(t)]$ with $t\in[2.5,7.5]$, as shown in FIG. 3 (b). The peaks are markedly enhanced with increasing $u$, and the subsequent enhanced revivals will be significantly more pronounced.
Figure 3: (a) The representative fidelity dynamics for the disordered Hamiltonian $\hat{H}_{\rm dis}$ in Eq. (6) at several values of $u$. At $u{=}0$, the scarring signatures swiftly collapse at moderate disorders. The Fock skin effect at up to $u{=}0.35$ leads to prolonged and robust scarring recurrences. (b) The first fidelity peak $F_{1}$, observed with $t\in[2.5,7.5]$, which is significantly shifted by non-Hermiticity $u$. These results in (b) are based on $200$ realizations, and the error bar denotes the standard deviation obtained throught such $200$ realizations. All the results in (a) and (b) are computed under conditions: system size $L{=}14$, |$\ket{\mathbb{Z}_{2}}$ initial state, and uniform disorder strength of $W=0.8$.
Optical lattice proposal.—Finally, we outline a proposal for observing non-Hermitian QMBSs in existing Bose-Hubbard (BH) quantum simulators with additional laser-induced loss [71, 72, 73, 32], see FIG. 4. Such a simulator is described by following BH-type Hamiltonian (see SM [66] for a derivation)
@@
where $\hat{b}_{j}^{\dagger}$ ($\hat{b}_{j}$) denotes the boson creation (annihilation) operator on site $j$, and $\hat{n}_{j}{=}\hat{b}^{\dagger}_{j}\hat{b}_{j}$. The nonreciprocity $J{\pm}\tilde{\gamma}$ results from the adiabatic elimination of a dissipative excited state (wavy arrow in FIG. 4 for laser-induced atom loss) linked to the nearest ground states [59]. The potential tilt $\Delta$ and on-site repulsion $U$ induce kinetic constraints. The effective non-Hermitian PXP model (1) arises when we set the boson filling factor to $\nu{=}N/L{=}1$ and $U{=}\Delta\gg J,\gamma$ [66]. Under these conditions, the local transition between the “20” doublon and the “11” neighboring single-boson configurations can be exactly mapped to the Pauli spin flip operator $\hat{X}$ [74, 75, 32, 66], as shown by the red box in Fig. 4 (a), such that the initial |$\ket{\mathbb{Z}_{2}}$ state corresponds to a product state of doublons separated by empty sites, i.e. “$2020$”.
@@
Enhanced scarring can be experimentally observed by measuring the total imbalance of the boson density across odd and even sites, $I(t){=}(1/N){\sum_{i}\left(n_{2i}(t){-}n_{2i+1}(t)\right)}$, where $n_{i}(t){=}\bra{\psi(t)}\hat{n}_{i}\ket{\psi(t)}/\bra{\psi(t)}\ket{\psi(t)}$ is the normalized occupation of site $i$ and $N$ is the total number of bosons. $I(t)$ is proportional to the Néel order parameter or |$\ket{\mathbb{Z}_{2}}$ echo corresponding to the effective spins. Simulation results are shown in FIG. 4(b) with $V{=}\Delta{=}10J$. Saliently, the non-Hermitian case with $\tilde{\gamma}{=}0.3J$ (red) exhibits stronger revivals of $I(t)$ than the Hermitian case with vanishing $\tilde{\gamma}$ (blue).

### v2

@@ Abstract
In contrast with extended Bloch waves, a single particle can become spatially localized due to the so-called skin effect originating from non-Hermitian pumping. Here we show that in kinetically-constrained many-body systems, the skin effect can instead manifest as dynamical amplification within the Fock space, beyond the intuitively expected and previously studied particle localization and clustering. We exemplify this non-Hermitian Fock skin effect in an asymmetric version of the PXP model and show that it gives rise to ergodicity-breaking eigenstates – the non-Hermitian analogs of quantum many-body scars. A distinguishing feature of these non-Hermitian scars is their enhanced robustness against external disorders. We propose an experimental realization of the non-Hermitian scar enhancement in a tilted Bose-Hubbard optical lattice with laser-induced loss. Additionally, we implement digital simulations of such scar enhancement on the IBM quantum processor. Our results show that the Fock skin effect provides a powerful tool for creating robust non-ergodic states in generic open quantum systems.
@@
The experimental observations of QMBSs in synthetic platforms ranging from ultracold atoms to superconducting circuits [11, 29, 30, 31, 32, 33, 34] have revealed the intrinsic fragility of QMBS dynamical revivals. In some cases, this has been mitigated by periodic driving, which can extend the lifetime of QMBS oscillations [29, 32, 35, 36, 37]. Nevertheless, thermalization or decay to a steady state, due to inherent decoherence from atom loss or thermal fluctuations, ultimately sets in. Thus, achieving stable QMBSs in real-world settings, with unavoidable non-unitarity, remains a challenge [38, 39, 40, 41, 42].
In this paper, we show that non-Hermiticity presents a new route for enhancing ergodicity breaking and many-body scarring. Borrowing inspiration from non-Hermitian pumping, which breaks the conventional bulk-edge correspondence [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], we demonstrate that a class of many-body systems with kinetic constraints admit a form of “Fock skin effect” that amplifies the revivals of special initial states and makes them resilient against spatial disorder. Unlike the conventional non-Hermitian skin effect (NHSE), which can induce nontrivial edge localization in real space, the Fock skin effect is characterized by persistent dynamical recurrences in the Fock space of a many-body system. We demonstrate this mechanism of QMBS enhancement in a non-Hermitian variant of the PXP model, which can be realized in a Bose–Hubbard optical lattice with laser-induced loss [57, 58, 59, 60, 61]. Connection with conventional non-Hermitian pumping is made through the “forward scattering approximation” [12, 62, 22], which effectively maps the many-body system to a spatially-inhomogeneous tight-binding chain with non-reciprocal hopping amplitudes. Such QMBS enhancement is further demonstrated in the digital simulation conducted on the IBM Q device without deliberate error mitigation. Our results offer a new perspective on obtaining stable many-body revivals through dissipative perturbations, extending beyond recent approaches based on decoherence-free subspaces [63, 64].
Figure 1: The directed Fock-space graph of the non-Hermitian PXP model in Eq. (1) with $L{=}6$ sites under PBCs. Each vertex represents a Fock state, organized into layers by the Hamming distance $x$ from the state |$\ket{\mathbb{Z}_{2}}$ on the left. The translated |¯$\ket{\bar{\mathbb{Z}}_{2}}$ state is on the opposite end. Edges represent the transitions allowed by the Hamiltonian in Eq. (1). Red links represent the symmetric coupling $\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$ (example of $P_{0}X_{1}P_{2}$ shown). Blue directed arrows label the asymmetric couplings $\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}$ (example of $P_{3}X_{4}^{\prime}P_{5}$ shown), with the arrow denoting enhanced spin flipping $(1{+}u)\hat{X}^{+}_{j}$. All directed edges point towards $x{=}0$, denoting asymmetric pumping from the Fock skin effect.
@@
where $\hat{X}$, $\hat{Y}$, $\hat{Z}$ denote the standard Pauli matrices, with $\hat{P}{=}(\mathds{1}{-}\hat{Z})/2$ projecting onto the spin-$\downarrow$ state, $j=0,1,\ldots,L-1$ and $L$ is the number of spins. Non-Hermiticity enters through the asymmetric spin flip operator, $\hat{X}^{\prime}{=}\hat{X}{+}iu\hat{Y}$, acting on the even sites, FIG. 1, where $u$ controls the strength of non-Hermiticity. Each $\hat{X}$ or $\hat{X}^{\prime}$ spin flip acts only when its neighboring spins are both spin-$\downarrow$. To distinguish our Fock skin effect from conventional skin effects with spatial boundary accumulation, we use periodic boundary conditions (PBCs). We also restrict $10{<}u{<}$ 1 to keep the energy spectrum real and avoid wavefunction decay. More results on the irrelevance of physical boundaries and level spacing distribution are given in Supplementary Material (SM) [67].
@@
To provide insight behind the non-Hermitian scar enhancement, we analyze the wavefunction dynamics within the Fock space. The latter can be visualized as a quasi-1D network structure by grouping Fock states according to their Hamming distance, which measures the length of the shortest path from any Fock state and the Néel state, $|\ket{\mathbb{Z}_{2}}=\ket{\uparrow\downarrow\uparrow\downarrow\cdots}$, see an example in FIG. 1. Note that the |$\ket{\mathbb{Z}_{2}}$ state is the one that exhibits QMBS revivals in the Hermitian PXP model [12]. The Hamming distance indicates how many applications of the Hamiltonian are needed to reach the states in a given layer starting from the |$\ket{\mathbb{Z}_{2}}$ state. This is because the Hamiltonian only connects layer $x$ with $x\pm 1$, as shown by edges in FIG. 1, making apparent the connection with a tight-binding chain. Crucially, this picture reveals a consistent direction in the asymmetry of the transitions between Fock states. As illustrated, the blue directional transitions are all directed towards the |$\ket{\mathbb{Z}_{2}}$ state at $x{=}0$, reminiscent of a conventional NHSE chain but with Fock states instead of physical sites. However, with multiple Fock states in each layer instead of a single site, the “boundary” skin localization would be weak [67], and we instead expect more pronounced QMBS revivals, rather than the directional bulk current in conventional NHSE systems [44, 47, 68, 69].
@@
The stratification of the Fock space into Hamming layers can be formally expressed by generalizing the forward scattering approximation (FSA) [12, 62], which allowed to understand the origin of QMBSs in the Hermitian case [22]. Here, we demonstrate that this method can approximate the behavior of scar states in our non-Hermitian setup. We first decompose our non-Hermitian PXP Hamiltonian in Eq. (1) as $\hat{H}{=}\hat{H}^{+}{+}\hat{H}^{-}$, where
$\hat{H}^{+}=$ $\sum_{j\in{\rm even}}\hat{P}_{j-1}\hat{X}^{\prime-}_{j}\hat{P}_{j+1}\!+\!\sum_{j\in{\rm odd}}\hat{P}_{j-1}\hat{X}^{+}_{j}\hat{P}_{j+1},$ (2)
$\hat{H}^{-}=$ $\sum_{j\in{\rm even}}\hat{P}_{j-1}\hat{X}^{\prime+}_{j}\hat{P}_{j+1}\!+\!\sum_{j\in{\rm odd}}\hat{P}_{j-1}\hat{X}^{-}_{j}\hat{P}_{j+1},$
@@
with $\hat{X}^{\pm}{\equiv}(\hat{X}\pm i\hat{Y})/2$ and $\hat{X}^{\prime\pm}{\equiv}(1\pm u)(\hat{X}\pm i\hat{Y})/2$. By successively acting with $\hat{H}^{+}$ on the |$\ket{\mathbb{Z}_{2}}$ state, which we label as |$\ket{0}_{\rm FSA}$, we can generate a tower of $L+1$ FSA basis states |$\ket{n}_{\mathrm{FSA}}{\propto}(H^{+})^{n}\ket{0}_{\mathrm{FSA}}$, with $n=0,1,2,\ldots,L$ [62]. These states span the Fock-space layers in FIG. 1. In the FSA basis, the projected Hamiltonian (1) is given by
@@
with the non-uniform and asymmetric hopping amplitudes $t^{+}_{x}=\bra{x\!+\!1}_{\rm FSA}\!\!\hat{H}^{+}\!\!\ket{x}_{\rm FSA}$ and $t^{-}_{x}=\bra{x}_{\rm FSA}\!\!\hat{H}^{-}\!\!\ket{x\!+\!1}_{\rm FSA}$. Despite a formal analogy between the model (3) and a tight-binding chain, each FSA basis state |$\ket{x}_{\mathrm{FSA}}$ represents a superposition of all Fock states in the Hamming layer $x$. Moreover, both the basis states and the effective hoppings $t^{\pm}$ intricately depend on the non-hermiticity $u$ of the Hamiltonian, as we show below.
@@
Figure 2: (a) Normalized fidelity $F(t)$ for the initial state |$\ket{\mathbb{Z}_{2}}$ in the non-Hermitian PXP model (1). Moderate non-Hermiticity $u\neq 0$ significantly enhances the return probability to the initial state. For large $u$ (dashed green curve), robust but irregular revivals occur due to multiple frequencies in the dynamics. (b) Amplitudes of effective hoppings $t_{x}^{\pm}$. The asymmetry between $t^{\pm}_{x}$ for $u{>}0$ (circles vs. triangles) leads to the non-Hermitian Fock skin effect in our model. (c) The concentration of eigenstates in the Hamming layer $x$, Eq. (4). A sharp Fock skin mode emerges under strong non-Hermiticity (green curve), corresponding to the persistent revivals in (a). (d) Comparison between the FSA left-boundary localization, $|\bra{\mathbb{Z}_{2}}\ket{\phi_{j}}_{\rm FSA}|$ |, and overlaps of |$\ket{\mathbb{Z}_{2}}$ with all eigenstates of the Hamiltonian in Eq.(1). From $u{=}0.0$ to $u{=}0.2$, the FSA boundary localization closely matches the overlap of $\mathbb{Z}_{2}$ state with the equally-spaced QMBS states. All data is for a chain of size $L{=}16$ with PBCs.
@@
QMBS enhancement due to Fock skin effect.—QMBSs manifest as pronounced revivals in the fidelity, $F(t)\equiv|{\bra{\psi(0)}\ket{\psi(t)}}|^{2}/{\bra{\psi(t)}\ket{\psi(t)}}$ with |$\ket{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, from the |$\ket{\psi(0)}=\ket{\mathbb{Z}_{2}}$ state for PXP-type models [11, 12]. Strong revivals in $F(t)$ indicate broken ergodicity since the fidelity should decay rapidly under thermalizing dynamics [12, 13]. In FIG. 2(a), the fidelity $F(t)$ is plotted for the PXP model (1) with various degrees of non-Hermiticity $u$. The decay of peaks in $F(t)$ is noticeably slower for $u{=}0.2$ (blue) compared to the Hermitian $u{=}0$ case (red). With much larger non-Hermiticity $u{=}0.7$ (green dashed), both frequency and magnitude of the revivals are significantly enhanced, showing no significant decay even after $t{=}10$.
@@
To explicitly attribute this non-Hermitian scar enhancement to the Fock skin effect, we plot the hopping amplitudes $t_{x}^{\pm}$ in FIG. 2(b). Saliently, the FSA hoppings are asymmetric, with the disparity between $t_{x}^{+}$ (round markers) and $t_{x}^{-}$ (triangle markers) becoming more pronounced with greater non-Hermiticity $u$. This asymmetry is markedly different from that in the physical Hamiltonian (1), since $x$ is the Hamming distance from the |$\ket{\mathbb{Z}_{2}}$ state in Fock space and not a spatial coordinate. Importantly, these hopping amplitudes are non-trivially dependent on $x$ but consistently asymmetric ($t_{x}^{+}\geq t_{x}^{-}$) along the FSA chain. This property is not generic, see SM [67] for the discussion of a related model where this is not the case, with crucial implications on the existence of Fock skin accumulation.
@@
As seen in FIG. 2(c), the Fock skin pumping of $u{=}0.2$ shifts the symmetric $\Pi_{\rm Fock}(x)$ profile of $u{=}0$ slightly leftward, yet without pronounced “boundary” localization. Nevertheless, this increased proximity to the |$\ket{\mathbb{Z}_{2}}$ state is sufficient to noticeably enhance the scarred revivals [FIG. 2(a), blue]. Stronger non-Hermiticity $u{=}0.7$ results in a more skewed $\Pi_{\rm Fock}(x)$, but less localized than conventional skin states with $t_{x}^{+}/t_{x}^{-}{\approx}3$ [FIG. 2 (b), green], with localization lengths around $[\log 3]^{-1}{\approx}1$ site. Crucially, this lack of conventional exponential boundary skin localization allows for significant state diffusion in the Hilbert space, but with enhanced return probability to the |$\ket{\mathbb{Z}_{2}}$ state.
@@
Figure 3: (a) Fidelity dynamics for the |$\ket{\mathbb{Z}_{2}}$ initial state evolving under the disordered Hamiltonian $\hat{H}_{\rm dis}$ at several values of $u$. At $u{=}0$, scarring signatures swiftly collapse at moderate disorders, while the Fock skin effect leads to prolonged and robust scarring recurrences up to $u{=}0.35$. (b) The first fidelity peak $F_{1}$ within $t\in[2.5,7.5]$, which is significantly shifted by non-Hermiticity $u$. Error bars represent the standard deviation over $200$ realizations. All data is for system size $L{=}14$ and uniform disorder strength of $W=0.8$.
Robustness to disorder.— We have established that the Fock skin effect can significantly enhance QMBSs. Given that the real-space skin effect can result in robust, directed dynamical currents, we posit that the non-Hermiticity in our system can not only amplify the quantum scar but also protect it against external (Hermitian) perturbations. To check this, we investigate the robustness of quantum scar under spatial disorder, $\hat{H}_{\rm dis}=\hat{H}+\sum_{j}w_{j}\hat{Z}_{j}$, where $w_{j}$ denotes the strength of on-site random potential drawn from a uniform distribution within the range $[-W/2,W/2]$. It was previously shown that Hermitian QMBS revivals decay with increasing disorder strength [71]. For our model with the Fock skin effect at up to $u{=}0.35$, this resilience under moderate disorders is remarkably highlighted by the fidelity shown in FIG. 3(a), where robust scarring signatures are still present, in contrast to the rapid collapse at Hermitian $u{=}0$ without the Fock skin effect. The enhanced stability is demonstrated through the first fidelity peak, $F_{1}={\rm Max}[F(t)]$ with $t\in[2.5,7.5]$, as shown in FIG. 3 (b). $F_{1}$ as well as subsequent peaks are markedly enhanced with increasing $u$.
Optical lattice proposal.—The Fock skin effect and non-Hermitian QMBSs can be observed in a Bose-Hubbard (BH) quantum simulator with additional laser-induced loss [72, 73, 74, 32]. The setup is illustrated in FIG. 4 (a) and it corresponds to the Hamiltonian
@@
where $\hat{b}_{j}^{\dagger}$ ($\hat{b}_{j}$) denotes the boson creation (annihilation) operator on site $j$, the occupation number $\hat{n}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$, and we use open boundary conditions (OBCs), see SM [67] for a derivation. The nonreciprocity $J\pm\tilde{\gamma}$ results from coupling a dissipative excited state (wavy arrow in FIG. 4 for laser-induced atom loss) to its nearest ground states [60].
The tilt $\Delta$ and on-site repulsion $U$ induce kinetic constraints under the condition $U=\Delta\gg J,\tilde{\gamma}$ and with the boson filling factor set to $\nu=N/L=1$, where $N$ is the total number of bosons [32, 67]. In this regime, the three-boson occupancy of any site is forbidden and the only allowed dynamical processes consist of local hops $\cdots 20\cdots\leftrightarrow\cdots 11\cdots$. The latter can be mapped to the Pauli $\hat{X}$ operator [75, 76], as shown by the red box in Fig. 4 (a). In the resulting effective non-Hermitian PXP model (1), the |$\ket{\mathbb{Z}_{2}}$ state simply maps to a product state of doublons separated by empty sites, |$\ket{2020\ldots}$.
@@
In FIG. 4(b) we numerically compute the dynamics of boson imbalance, $I(t)=(1/N){\sum_{i}\left(n_{2i}(t){-}n_{2i+1}(t)\right)}$, when the system is prepared in the |$\ket{2020\ldots}$ state, with the normalized occupations $n_{i}(t)=\bra{\psi(t)}\hat{n}_{i}\ket{\psi(t)}/\bra{\psi(t)}\ket{\psi(t)}$. $I(t)$ is proportional to the |$\ket{\mathbb{Z}_{2}}$ echo of the effective spins. The results in FIG. 4(b) with $U=\Delta=10J$ show that the non-Hermitian case with $\tilde{\gamma}=0.3J$ (red) exhibits stronger revivals of $I(t)$ than the Hermitian case with vanishing $\tilde{\gamma}$ (blue).